# The Geometric View of Theories

### v1

@@ Abstract
Recent critiques of the semantic conception of scientific theories suggest that a theory is not best formulated as a collection of models satisfying some set of kinematical or dynamical conditions. Thus it has been argued that additional structure on the set of models is required. Furthermore, there are calls for developing a ‘theory of theories’, where what was formerly a ‘theory’ is seen as a ‘model’ within a larger theoretical structure. This paper makes a two-pronged proposal for the “shape” that physical theories should take, based on recent insights on dualities and quasi-dualities in physics. First, I develop a geometric view of theories, according to which a physical theory is a set of models with topological and geometric structure on it. This general view is briefly illustrated in an example from quantum cosmology. Second, I make a more specific proposal for a natural structure that can encompass various ‘theories’ as its models, with topological and algebraic-geometric structure on them. I call the latter more specific structure a ‘model bundle’, where the models are in the fibres and there is a moduli space in the base. I illustrate my second proposal in the Seiberg-Witten theory (Seiberg and Witten, 1994a,b), where the moduli space is the complex plane with three punctures, the states and quantities are in the fibres, and the modular group is the structure group that acts on the fibres. This view highlights the important role of quasi-dualities as local transition functions between fibres; dualities are recovered as global transition functions when the bundle is trivial. I discuss some philosophical issues that this geometric view of physical theories opens up, such as its realist interpretation.
@@ 1 Introduction
Recent discussions have emphasised the kinship between these two views: models were always integral to the syntactic view, and likewise the semantic view was—despite some radical-sounding initial claims by its proponents—never meant to dispense with language. In fact, one speaks of a recent consensus, according to which both aspects are needed. I will call this consensus view the ‘combined view’: [^1] See Halvorson (2013), Glymour (2013), van Fraassen (2014: p. 276), and Lutz (2017: p. 330). An overview of the subject is given in Frigg (2023), who discusses the consensus at p. 167. However, there is also dissent: Wallace’s (2022: p. 349; 2024: pp. 1-2) distinction between ‘language-first’ and ‘maths-first’ approaches to physical theories widens the gap between the syntactic and semantic conceptions. This distinction seems to be chiefly motivated by the aim to develop the formal underpinnings of a structural realist metaphysics. namely, the view that a scientific theory requires both a language and a set of models, where the language need not be first-order logic. On this view, the linguistic aspect includes a commitment to a set of sentences, and each model makes these sentences true.
@@
(i)  Inter-theoretic relations: These suggest that there are formal and structural relationships between models that the standard conceptions of theories usually treat as unrelated. Thus discussions of equivalence of theories emphasise different ways in which models or theories can be equivalent or inequivalent. [^2] For an overview of these developments, see Weatherall (2019). Likewise, relations of reduction and correspondence suggest the need to develop formal notions of ‘closeness’ and ‘membership’ of theories, so as to have criteria for when a theory approximates, and when it subsumes, another theory. Therefore, to go beyond the combined view, one needs to choose the inter-theoretic relations that one wishes to include as part of one’s conception of ‘theory’. Consquently, these inter-theoretic relations, especially equivalence relations, may prompt a change in one’s usage of ‘theory’ and ‘model’.
@@
Recent proponents of the structured view of theories include Halvorson (2012: pp. 204–205; 2019: p. 277) and Halvorson and Tsementzis (2017: pp. 413–414). [^4] See also Curiel (2014: p. 275; cf. p. 318) and North (2021: pp. 187-188). In general relativity, spaces of solutions of the Einstein field equations, and criteria for closeness, were discussed very early on (see Hawking and Ellis, 1973: pp. 249, 252-254). For a philosophical discussion, see Belot (2018: p. 968); also Fletcher (2016: p. 366). For example, Geroch (1969: p. 182) constructed a five-manifold from a one-parameter family of four-manifolds that depend smoothly on a continuous parameter. His motivation was to define the limit of this sequence of four-dimensional metrics, as the metric on the boundary of the five-dimensional manifold. They motivate their requirement for having ‘relevant topological information’ on the set of models by the question of when a theory, understood as a set of sentences, can be reconstructed from its category of models. Lehmkuhl (2017: pp. 1-10) is a programmatic call for a ‘theory of theories’. The first (modest) step in this programme is to compare individual theories.
@@
I will argue that there are two key notions from these areas of physics that need to be incorporated into the philosophical discussion of the structure of physical theories: namely, dualities i.e. isomorphisms of theories, and their generalizations which I call quasi-dualities (where the map falls short of being an isomorphism), together with the idea of a moduli space and, more generally, a space of parameters. Although spaces of theories and moduli spaces may be familiar from quantum gravity and other areas of physics and mathematics: exceptions aside, virtually no trace of these notions can be found in the philosophical literature.Similarly, the role of parameter spaces in scientific theories is seldom discussed. This stands in stark contrast to the contemporary physics literature, where, for good reasons that I hope to illustrate, these notions take centre stage.
@@
(1)  Lifting ‘theory’ and ‘model’ “one level up”. Advances in physics, particularly the discovery of dualities, prompt us to be flexible about the use of the words ‘theory’ and ‘model’. Specifically, we will be primarily concerned with how to formulate a ‘theory of theories’. What was formerly called ‘theory’ is now called ‘model’ so that a model will be (a) general and not a specific solution and (b) exact and not approximate—so lacking common connotations of ‘model’ in the (vast!) modelling literature. Thus according to the idea of generality discussed above, we lift our usage of ‘theory’ and ‘model’ “one level up”. [^5] De Haro (2016: p. 261) and De Haro and Butterfield (2018: p. 309) already proposed that dualities prompt moving one’s usage of ‘model’ and ‘theory’ one level up. In this paper, I defend this idea in contexts where the models are not duals, and also combine it with the second thesis, namely that there is additional structure on the set of models that is now the theory, one level “up”.
@@
Contrast this with the literature on theoretical equivalence cited above, which typically focusses on formal criteria for comparing different theories, rather than expanding the notion of a ‘theory’ to encompass former theories as its models. Furthermore, while dualities are the main reason for the “up”-move in physics, dualities are not often considered in discussions of theoretical equivalence. [^6] A commendable exception is Weatherall (2016, 2020). This is of course not to deny that theoretical equivalence has been a main topic of discussion in the recent philosophical literature on dualities: for an overview of this literature, see De Haro and Butterfield (2025: pp. 93-98) and De Haro (2021). Hence my earlier qualification that the literature on theoretical equivalence addresses point (i) only ‘in part’.
@@
(2)  Structure on the set of models. My proposal will be that one should not view a theory as a collection of models, but rather as a single geometric object. Thus one equips one’s set of models successively with topological, differential, algebraic and geometric structure. For example, there is a norm from which one can define a metric: and often also other geometric structure, such as complex and symplectic structure. Since one wishes such structure to be physically significant, it is essential to take it from contemporary physics. It is the absence of such algebraic-geometric structures in discussions of theoretical equivalence, together with additional structures that the next Section will review in detail, that prompts my previous qualification: the literature addresses point (ii) ‘not at all’. [^7] Concrete proposals for equipping the set of theories with structure include those by Halvorson and Tsementzis (2017), mentioned earlier. They propose to discuss the “universe” of scientific theories as a two-category of categories. In addition, topological structure is discussed in Halvorson (2019: p. 79ff). However, it is fair to say that, on balance, their work remains in the realm of logic, and does not consider structures of the algebraic-geometric type that I will discuss in this paper, which are required to describe contemporary physical theories. This comment is bolstered by their stated ambition to describe a “universe” of theories. For in the complex realm of physical theories, it is best to start by describing a single “planetary system” in detail before attempting to describe a whole “universe”. This is not a criticism, but a matter of their prioritizing “toy” theories expressed in a formal logical language.
@@ 2.1 Dualities: lifting the usage of ‘theory’ “one level up”
A duality is an isomorphism between physical theories. This isomorphism is defined with respect to the structure that the duals share. If this structure is itself a bare theory, then the isomorphism is a duality. By a ‘bare theory’, I mean basic structure that comprises a physical theory, regardless of interpretation: here, I follow De Haro and Butterfield (2018) in taking this structure to be a set of states, a set (almost always: an algebra) of quantities, and a dynamics, together with a rule for assigning values to quantities on states. [^10] De Haro and Butterfield (2025) stress that this is one of the “shapes” that a bare theory or model can take. The dynamics can be either deterministic, as in classical mechanics, or probabilistic, as in statistical mechanics. Also, there are alternatives, in terms of path integrals, to the formulation in terms of states, quantities and dynamics. For a discussion, see De Haro, Teh and Butterfield (2017: p. 75). (We will return to this definition in the next Section.) This shared structure is called the common core. The structure that duals do not share is called specific structure.
@@
Although this choice for what we call ‘theory’ and ‘model’ may initially appear to be a mere choice of words: if combined with point (2), about structure, it is substantive. [^11] De Haro (2016: pp. 266–270) already discussed that the choice is substantive once it is combined with an interpretation, which can be internal or external. For it will lead us to a new formulation of the type of structure that we use to formulate scientific theories, and to a cautious realism about this structure (see Section 3).
@@ 2.2 Structure on the set of models: quasi-dualities and moduli spaces
A quasi-duality is a map between models that falls short of being a duality, i.e. it falls short of being an isomorphism that preserves a common core theory. This can happen in either of the two ways in which a map can fall short of satisfying this definition: namely, through its failing to be a bijection, and through its failing to preserve a common core that is a theory. For example, it sometimes happens that a quasi-duality preserves some structure that the models have in common, but not enough, so that this common structure does not amount to a theory, i.e. a triple $T=\langle{\cal S},{\cal Q},{\cal D}\rangle$ of states or state-space $\cal S$}, algebra of quantities $\cal Q$} and dynamics $\cal D$}, together with a set of values for the quantities that also match between the duals.
@@
As a second example, take another space of metrics: namely, the space of Einstein metrics on a compact smooth manifold up to the diffeomorphisms of the manifold that are continuously connected to the identity. Here, an Einstein metric is one whose Ricci tensor is, up to a constant, equal to the metric tensor (and we consider the case that the metric tensor has positive signature, i.e. it is a Riemannian manifold). This constant can be positive, zero, or negative, where the constant being zero is the familiar case of vacuum general relativity. These classes of isometric Riemannian manifolds of constant curvature, also called ‘Einstein structures’, form a finite-dimensional moduli space (Besse, 1980: pp. 340, 358). In general, for a given manifold, its moduli space of Einstein structures is not a manifold (for example, some of these moduli spaces have singular points, or are disconnected), although it does have a covering that is a smooth manifold. [^23] For some of the properties of the moduli space, like its having constant scalar curvature, its being Hausdorff, and its having isolated points that cannot be deformed in the way discussed above, so that it corresponds to a unique object that does not belong to a continuous family of related objects, see Besse (1980: pp. 352-355). Examples of moduli spaces and local deformations are at pp. 365-366; see also Kodaira (1986: pp. Chapter 4).
@@
For example, the metric of the moduli space, which is in general different in an obvious way from a spacetime metric, lives on the tensor bundle of the moduli space, i.e. it assigns a symmetric bilinear form on each tangent space. And, as we will see in the examples in Appendix A, the models, each comprising a set of states and quantities, live on fibres over the moduli space. The main idea of fibre is a class of objects—for us: models—associated with (so imagined as “standing above’) an element of the base-space—for us: the moduli space. Agreed, in the context of differential geometry, ‘fibre bundle’ has a technical geometric meaning that includes e.g. smoothness conditions. I will indeed be committed to this technical meaning, which recurs in my main example in Appendix A. However, in other examples that are formulated using algebraic geometry, the base-space can be an algebraic variety. The analogous (more general) structure required is then a sheaf, which can be defined using algebraic properties, rather than smoothness. I will briefly return to such generalizations in Section 3.1.3.
@@
In view of the above discussion, it is natural to propose that in general, for a large class of interesting examples, we have a bundle whose sections are the values of quantities, varying over the moduli space. The central idea that leads to the geometric view of theories is to apply the notion of a moduli space as a space of parameters of some family of objects, to scientific theories. We do this by taking the objects to be models. This will be my main proposal in Section 3: I will call the resulting bundle, with a moduli space in the base and a model (states and quantities) in the fibres, a model bundle.
@@
Note that model bundles are very different to the one obvious place where fibre bundles appear in physics: namely, principal bundles as the mathematical backbone of gauge theories. Principal bundles are used in the context of classical field theory to incorporate local gauge symmetries of fields. By putting the structure group in the fibre, the group elements in the fibre are allowed to vary over the spacetime, which gives rise to a larger group of local gauge transformations, namely the group of bundle automorphisms. In this way, principal bundles are helpful in making the redundancies of gauge theories manifest. By contrast, in a model bundle there are models in the fibre, i.e. sets of states and quantities that have a physical interpretation, typically characterized by vector bundles: [^24] For more on how vector bundles can characterize states and quantities, see Appendix A.4. and there is a moduli space of parameters or fields in the base manifold. The structure group is the group of quasi-dualities, which are not redundancies but rather relate distinct physical situations. The contrast can be summed up as follows: principal bundles give us a way to handle redundancy in gauge theory, while model bundles give us a way to handle physical states and quantities that depend on parameters.
@@
In this bundle picture, moduli spaces will always be either manifolds, perhaps with singularities, or algebraic varieties (and then one expects to generalize the bundle to a sheaf). The whole bundle need not be a manifold or an algebraic variety. Also, we will see that a model bundle covers a large class of cases where the fibres are isomorphic to each other. If examples exist where the fibres are not isomorphic, these require a further generalization (see the discussion in Section 3.1.3). In other words, my strategy is to build my way up from examples in physics, using the minimal mathematical level of generality required, and generalizing only as required by the examples, rather than on a priori mathematical grounds. Using an astronomical metaphor, my aim is to start by describing theories that are familiar “planetary systems”, before attempting to describe the whole “universe” of possible theories (cf. footnote 7).

### v2

@@ Abstract
Recent critiques of the semantic conception of scientific theories suggest that a theory is not best formulated as a collection of models satisfying some set of kinematical or dynamical conditions. Thus it has been argued that additional structure on the set of models is required. Furthermore, there are calls for developing a ‘theory of theories’, where what was formerly a ‘theory’ is seen as a ‘model’ within a larger theoretical structure. This paper makes a two-pronged proposal for the “shape” that physical theories should take, based on recent insights on dualities and quasi-dualities in physics. First, I develop a geometric view of theories, according to which a physical theory is a set of models with topological and geometric structure on it. This general view is briefly illustrated in an example from quantum cosmology. Second, I make a more specific proposal for a natural structure that can encompass various ‘theories’ as its models, with topological and algebraic-geometric structure on them. I call the latter more specific structure a ‘model bundle’, where the models are in the fibres and there is a moduli space in the base. I illustrate my second proposal in an example from quantum field theory. This view highlights the important role of quasi-dualities as local transition functions between fibres; dualities are recovered as global transition functions when the bundle is trivial. I discuss some philosophical issues that this geometric view of physical theories opens up, such as its realist interpretation.
Forthcoming in Synthese.
@@ 1 Introduction
Recent discussions have emphasised the kinship between these two views: models were always integral to the syntactic view, and likewise the semantic view was—despite some radical-sounding initial claims by its proponents—never meant to dispense with language. In fact, one speaks of a recent consensus, according to which both aspects are needed. I will call this consensus view the ‘combined view’: [^1] See Halvorson (2013), Glymour (2013), van Fraassen (2014: p. 276), and Lutz (2017: p. 330). An overview of the subject is given in Frigg (2023), who discusses the consensus at p. 167. However, there is also dissent: Wallace’s (2022: p. 349; 2024: pp. 1–2) distinction between ‘language-first’ and ‘maths-first’ approaches to physical theories widens the gap between the syntactic and semantic conceptions. This distinction seems to be chiefly motivated by the aim to develop the formal underpinnings of a structural realist metaphysics. namely, the view that a scientific theory requires both a language and a set of models, where the language need not be first-order logic. On this view, the linguistic aspect includes a commitment to a set of sentences, and each model makes these sentences true.
@@
(i)  Inter-theoretic relations: These suggest that there are formal and structural relationships between models that the standard conceptions of theories usually treat as unrelated. Thus discussions of equivalence of theories emphasise different ways in which models or theories can be equivalent or inequivalent. [^2] For an overview of these developments, see Weatherall (2019). Likewise, relations of reduction and correspondence suggest the need to develop formal notions of ‘closeness’ and ‘membership’ of theories, so as to have criteria for when a theory approximates, and when it subsumes, another theory. Therefore, to go beyond the combined view, one needs to choose the inter-theoretic relations that one wishes to include as part of one’s conception of ‘theory’. Consequently, these inter-theoretic relations, especially equivalence relations, may prompt a change in one’s usage of ‘theory’ and ‘model’.
@@
Recent proponents of the structured view of theories include Halvorson (2012: pp. 204–205; 2019: p. 277) and Halvorson and Tsementzis (2017: pp. 413–414). [^4] See also Curiel (2014: p. 275; cf. p. 318) and North (2021: pp. 187–188). In general relativity, spaces of solutions of the Einstein field equations, and criteria for closeness, were discussed very early on (see Hawking and Ellis, 1973: pp. 249, 252–254). For a philosophical discussion, see Belot (2018: p. 968); also Fletcher (2016: p. 366). For example, Geroch (1969: p. 182) constructed a five-manifold from a one-parameter family of four-manifolds that depend smoothly on a continuous parameter. His motivation was to define the limit of this sequence of four-dimensional metrics, as the metric on the boundary of the five-dimensional manifold. They motivate their requirement for having ‘relevant topological information’ on the set of models by the question of when a theory, understood as a set of sentences, can be reconstructed from its category of models. Lehmkuhl (2017: pp. 1–10) is a programmatic call for a ‘theory of theories’. The first (modest) step in this programme is to compare individual theories.
@@
I will argue that there are two key notions from these areas of physics that need to be incorporated into the philosophical discussion of the structure of physical theories: namely, dualities i.e. isomorphisms of theories, and their generalizations which I call quasi-dualities (where the map falls short of being an isomorphism), together with the idea of a moduli space and, more generally, a space of parameters. Although spaces of theories and moduli spaces may be familiar from quantum gravity and other areas of physics and mathematics: exceptions aside, virtually no trace of these notions can be found in the philosophical literature. [^5] The only in-depth philosophical discussions of the moduli space of a physical theory that I am aware of are De Haro and Butterfield (2025), on which the current paper builds, and Vistarini (2019). Vistarini discusses moduli spaces in string compactifications, for other purposes than the ones we pursue in this paper. Similarly, the role of parameter spaces in scientific theories is seldom discussed. This stands in stark contrast to the contemporary physics literature, where, for good reasons that I hope to illustrate, these notions take centre stage.
@@
(1)  Lifting ‘theory’ and ‘model’ “one level up”. Advances in physics, particularly the discovery of dualities, prompt us to be flexible about the use of the words ‘theory’ and ‘model’. Specifically, we will be primarily concerned with how to formulate a ‘theory of theories’. What was formerly called ‘theory’ is now called ‘model’ so that a model will be (a) general and not a specific solution and (b) exact and not approximate—so lacking common connotations of ‘model’ in the (vast!) modelling literature. Thus according to the idea of generality discussed above, we lift our usage of ‘theory’ and ‘model’ “one level up”. [^6] De Haro (2017: p. 261) and De Haro and Butterfield (2018: p. 309) already proposed that dualities prompt moving one’s usage of ‘model’ and ‘theory’ one level up. In this paper, I defend this idea in contexts where the models are not duals, and also combine it with the second thesis, namely that there is additional structure on the set of models that is now the theory, one level “up”.
@@
Contrast this with the literature on theoretical equivalence cited above, which typically focusses on formal criteria for comparing different theories, rather than expanding the notion of a ‘theory’ to encompass former theories as its models. Furthermore, while dualities are the main reason for the “up”-move in physics, dualities are not often considered in discussions of theoretical equivalence. [^7] A commendable exception is Weatherall (2016, 2020). This is of course not to deny that theoretical equivalence has been a main topic of discussion in the recent philosophical literature on dualities: for an overview of this literature, see De Haro and Butterfield (2025: pp. 93–98) and De Haro (2021). Hence my earlier qualification that the literature on theoretical equivalence addresses point (i) only ‘in part’.
@@
(2)  Structure on the set of models. My proposal will be that one should not view a theory as a collection of models, but rather as a single geometric object. Thus one equips one’s set of models successively with topological, differential, algebraic and geometric structure.888‘Successively’ here marks a conceptual and mathematical procedure: starting from a set of models, one adds, in order of dependency, topological, differential, algebraic, and geometric structure. This mirrors standard mathematical practice (and the chain of forgetful functors from geometry down to set theory) and is intended to make explicit the kinds of structure that occur in contemporary physics. The ordering does not reflect a commitment to ontological priority of e.g. sets over topology. For example, there is a norm from which one can define a metric: and often also other geometric structure, such as complex and symplectic structure. Since one wishes such structure to admit a physical interpretation, it is natural to take it from contemporary physics rather than from purely logical or mathematical considerations. It is the absence of such algebraic-geometric structures in discussions of theoretical equivalence, together with additional structures that the next Section will review in detail, that prompts my previous qualification: the literature addresses point (ii) ‘not at all’. [^9] Concrete proposals for equipping the set of theories with structure include those by Halvorson and Tsementzis (2017), mentioned earlier. They propose to discuss the “universe” of scientific theories as a two-category of categories. In addition, topological structure is discussed in Halvorson (2019: p. 79ff). However, it is fair to say that, on balance, their work remains in the realm of logic, and does not consider structures of the algebraic-geometric type that I will discuss in this paper, which are required to describe contemporary physical theories. This comment is bolstered by their stated ambition to describe a “universe” of theories. For in the complex realm of physical theories, it seems best to start by describing a single “planetary system” in detail before attempting to describe a whole “universe”. This is not a criticism, but a matter of their prioritizing “toy” theories expressed in a formal logical language.
@@ 2.1 Dualities: lifting the usage of ‘theory’ “one level up”
A duality is an isomorphism between physical theories. This isomorphism is defined with respect to the structure that the duals share: a duality obtains only when that common structure is itself a bare theory (so that not any shared structure yields a duality). By a ‘bare theory’, I mean basic structure that comprises a physical theory, regardless of interpretation: here, I follow De Haro and Butterfield (2018) in taking this structure to be a set of states or state-space $\cal S$}, a set (almost always: an algebra) of quantities $\cal Q$}, and a dynamics $\cal D$}, together with a rule for assigning values to quantities on states. Thus a bare theory is a triple $T=\langle{\cal S},{\cal Q},{\cal D}\rangle$, and under a duality the set of values for the quantities also match between the duals. [^12] The state-space and the set of quantities almost always come equipped with structures, for example symmetries. The duality then also needs to preserve these structures. For a discussion, see De Haro and Butterfield (2025: pp. 7, 52, 60). Also, De Haro and Butterfield stress that the triple as here discussed is one of the “shapes” that a bare theory or model can take. The dynamics can be either deterministic, as in classical mechanics, or probabilistic, as in statistical mechanics. There are alternatives, in terms of path integrals, to the formulation in terms of states, quantities and dynamics. For a discussion, see De Haro, Teh and Butterfield (2017: p. 75). This shared structure $T=\langle{\cal S},{\cal Q},{\cal D}\rangle$, together with its assignments of values, is called the common core. The structure that duals do not share is called specific structure. [^13] The distinction between common core and specific structure is a formal i.e. non-interpretative one, because the notion of duality is formal i.e. non-interpretative. Thus ‘specific’ does not have the connotation of ‘redundant’. For example, there are cases where the specific structure should be interpreted as being non-redundant, such as the temperature in the Kramers–Wannier duality of the Ising model. For a discussion of this, and the corresponding distinction between external and internal interpretations, see De Haro and Butterfield (2025: pp. 50–51) and De Haro (2017: pp. 261–264).
@@
Although this choice for what we call ‘theory’ and ‘model’ may initially appear to be a mere choice of words: if combined with point (2), about structure, it is substantive. [^14] De Haro (2017: pp. 266–270) already discussed that the choice is substantive once it is combined with an interpretation, which can be internal or external. For it will lead us to a new formulation of the type of structure that we use to formulate scientific theories, and to a cautious realism about this structure (see Section 3).
@@ 2.2 Structure on the set of models: quasi-dualities and moduli spaces
A quasi-duality is a map between models that falls short of being a duality, i.e. it falls short of being an isomorphism that preserves a common core theory. This can happen in either of the two ways in which a map can fall short of satisfying this definition: namely, through its failing to be a bijection, and through its failing to preserve a common core that is a theory. For example, it sometimes happens that a quasi-duality preserves some structure that the models have in common, but not enough, so that this common structure does not amount to a theory, i.e. a triple $T=\langle{\cal S},{\cal Q},{\cal D}\rangle$.
@@
As a second example, take another space of metrics: namely, the space of Einstein metrics on a compact smooth manifold up to the diffeomorphisms of the manifold that are continuously connected to the identity. Here, an Einstein metric is one whose Ricci tensor is, up to a constant, equal to the metric tensor (and we consider the case that the metric tensor has positive signature, i.e. it is a Riemannian manifold). This constant can be positive, zero, or negative, where the constant being zero is the familiar case of vacuum general relativity. These classes of isometric Riemannian manifolds of constant curvature, also called ‘Einstein structures’, form a finite-dimensional moduli space (Besse, 1980: pp. 340, 358). In general, for a given manifold, its moduli space of Einstein structures is not a manifold (for example, some of these moduli spaces have singular points, or are disconnected), although it does have a covering that is a smooth manifold. [^26] For some of the properties of the moduli space, like its having constant scalar curvature, its being Hausdorff, and its having isolated points that cannot be deformed in the way discussed above, so that it corresponds to a unique object that does not belong to a continuous family of related objects, see Besse (1980: pp. 352–355). Examples of moduli spaces and local deformations are at pp. 365–366; see also Kodaira (1986: pp. Chapter 4).
@@
For example, the metric of the moduli space, which is in general different in an obvious way from a spacetime metric, lives on the tensor bundle of the moduli space, i.e. it assigns a symmetric bilinear form on each tangent space. And, as we will see in the examples in Appendix A, the models, each comprising a set of states and quantities, live on fibres over the moduli space. The main idea of fibre is a family of objects—for us: models—associated with (so imagined as “standing above”) an element of the base-space—for us: the moduli space. Agreed, in the context of differential geometry, ‘fibre bundle’ has a technical geometric meaning that includes e.g. smoothness conditions. I will indeed be committed to this technical meaning, which recurs in my main example in Appendix A. However, in other examples that are formulated using algebraic geometry, the base-space can be an algebraic variety. The analogous (more general) structure required is then a sheaf, which can be defined using algebraic properties, rather than smoothness. The model bundle is not a locally trivial fibre bundle also in cases where around special parameter values (e.g. $\hbar=0$) there is no neighbourhood that admits a trivialization, and the fibre there is not isomorphic to a generic fibre (see footnote 33 for an example.) I will briefly return to such generalizations in Section 3.1.3.
@@
In view of the above discussion, it is natural to propose that in general, for a large class of interesting examples, we have a bundle whose sections are the values of quantities, varying over the moduli space. The central idea that leads to the geometric view of theories is to apply the notion of a moduli space as a space of parameters of some class of objects, to scientific theories. We do this by taking the objects to be models. This will be my main proposal in Section 3: I will call the resulting bundle, with a moduli space in the base and a model (states and quantities) in the fibres, a model bundle.
@@
Note that model bundles are very different to the one obvious place where fibre bundles appear in physics: namely, principal bundles as the mathematical backbone of gauge theories. [^27] There are different conventions in the literature about the use of the words ‘fibre bundle’ and ‘principal bundle’, which however will not cause any confusion. Physics-oriented texts usually take the typical fibre of a principal bundle to be a Lie group (Isham 1999, p. 220; Nakahara, 2003, p. 363), while mathematically-oriented texts often characterise principal bundles via the group-action (Hamilton, 2017, p. 207). Also, physics-oriented texts often define a fibre bundle as having a structure group (Nakahara, 2003, p. 350), while other texts do not require this explicitly. Nothing hinges on these conventions, and I will adopt the physics convention of equipping a fibre bundle with a structure group. Principal bundles are used in the context of classical field theory to incorporate local gauge symmetries of fields. By putting the structure group in the fibre, the group elements in the fibre are allowed to vary over the spacetime, which gives rise to a larger group of local gauge transformations, namely the group of bundle automorphisms. In this way, principal bundles are helpful in making the redundancies of gauge theories manifest. By contrast, in a model bundle there are models in the fibre, i.e. sets of states and quantities that have a physical interpretation, typically characterized by vector bundles: [^28] For more on how vector bundles can characterize states and quantities, see Appendix A.4. and there is a moduli space of parameters or fields in the base manifold. The structure group is the group of quasi-dualities, which in general are not redundancies but rather relate distinct physical situations. The contrast can be summed up as follows: principal bundles give us a way to handle redundancy in gauge theory, while model bundles give us a way to handle physical states and quantities that depend on parameters.
@@
In this bundle picture, moduli spaces will always be either manifolds, perhaps with singularities, or algebraic varieties (and then one expects to generalize the bundle to a sheaf). The whole bundle need not be a manifold or an algebraic variety. Also, we will see that a model bundle covers a large class of cases where the fibres are isomorphic to each other. In examples where the fibres are not isomorphic, these require a further generalization (see the discussion in Section 3.1.3). In other words, my strategy is to build my way up from examples in physics, using the minimal mathematical level of generality required, and generalizing only as required by the examples, rather than on a priori mathematical grounds. Using an astronomical metaphor, my aim is to start by describing theories that are familiar “planetary systems”, before attempting to describe the whole “universe” of possible theories (cf. footnote 9).